At 8:15 Bonita left her house to the airport at an average speed of R65km/h. At 9:00 her boyfriend Bernard left Bonita's house and followed her path at an average speed of 104km/h. Bonita and Bernard arrived at the airport at exactly the same time. A linear relationship exist between the time and the distance travelled. The time is dependent on the number of kilometers travelled. Determine the time they arrived at the airport


1
Expert's answer
2021-05-24T19:16:24-0400

Let the time Bonita needs to reach airport will be x hours.

Then, the time Bernard needs to reach airport will be "x-\frac{45}{60}=x-0.75" hours.

"65x=104(x-0.75)."

"65x=104x-78."

"x=\frac{78}{104-65}=2" hours.

The time they arrived at the airport is 8:15+2=10:15.